# Novel Long Distance Free Space Quantum Secure Direct Communication for Web 3.0 Networksthanks:

### v4

@@ I Introduction
In the evolution of satellite-based QKD, significant strides have been achieved since the launch of the Micius satellite in 2017, a pioneering endeavor that demonstrated the transmission of secure quantum keys over intercontinental distances[3]. Despite these advancements, satellite-based QKD faces persistent challenges, including substantial signal loss, atmospheric interference, and the need for precise alignment and tracking of satellite signals[4, 5, 6].
Challenges with QKD for Global Scale Communication: QKD systems experience significant signal loss and attenuation over long distances, limiting practical distances without quantum repeaters to a few hundred kilometers. Atmospheric disturbances further exacerbate signal degradation in free-space QKD, making it highly susceptible to environmental factors. For instance, substantial signal loss and atmospheric interference hinder the efficiency of QKD in long-distance communications[7, 8].
Extending QKD reach beyond terrestrial limits with satellites introduces additional challenges, including precise satellite alignment and significant signal loss during transmission. Implementing a global QKD network requires substantial infrastructure, including numerous satellites and ground stations, and the integration of quantum repeaters. The high cost and complexity of this infrastructure pose barriers to widespread adoption[9, 10].
Advantages of QSDC for Global Scale Communication In contrast, QSDC protocols, particularly our solution LF QSDC, offer more practical solutions for global communication. QSDC allows direct transmission of secure messages without prior key exchange, reducing communication complexity and eliminating intermediate key management. Additionally, QSDC protocols incorporate eavesdropping detection through quantum state disturbance, ensuring real-time detection of interference[11, 12]. Technologies like Quantum-Aware Low-Density Parity-Check (LDPC) coding, Pointing, Acquisition, and Tracking (PAT) systems, and Atmospheric Quantum Correction Algorithms (AQCA) improve LF QSDC reliability over long distances. These technologies address key challenges in satellite-based free-space communication, ensuring precise alignment and minimizing signal loss due to atmospheric disturbances. LF QSDC’s reliance on advanced error correction and adaptive technologies reduces the need for extensive infrastructure compared to QKD, making it more scalable and cost-effective. By leveraging existing satellite and communication infrastructure with minimal modifications, LF QSDC can be integrated seamlessly into current systems[13, 14].
@@
Recent advancements in QSDC have shown significant theoretical and experimental progress, enhancing the feasibility of direct quantum communications[16, 17, 18, 19, 20, 21]. Noteworthy developments include the device-independent approach against collective attacks, which has become pivotal for enhancing communication security [22]. Furthermore, the realization of QSDC over 100 km using time-bin and phase quantum states underlines the scalability of these technologies[23].
@@
The memory-free DL04 protocol that FL QSDC employs reduces the complexity of quantum state storage and management, making the system more practical and robust for real-world applications [15] as shown in Table I. Our novel 1uantum-aware LDPC coding scheme, PAT technologies, and AQCA mitigate issues related to atmospheric turbulence and alignment errors, which are prominent in satellite-based free-space quantum communication [24, 25, 26, 27].
@@ II-B Measurement Protocols and Two-way Communication
After measurement, Bob announces his choice of $\theta$ and the results of his measurements. Alice and Bob then jointly analyze the outcomes to check if $\theta$ aligns with $a_{2}$ ($a_{2}\oplus\theta=0$ or $1$), ensuring the correct basis was used and the transmission was secure. This process embodies the two-way communication required by the DL04 protocol, where feedback from Bob influences potential retransmissions by Alice or adjustments in the encoding strategy to correct any discrepancies or to enhance security.
@@ III-A3 Security Enhancement Techniques
  • The estimation can be formulated as solving a set of linear inequalities derived from decoy state intensities and detection rates [26].
@@ III-C1 Design
The AQCA consists of four main modules: a disturbance modeler, a quantum error corrector, an adaptive optics system, and a signal enhancer and recoverer. The disturbance modeler is responsible for estimating the atmospheric parameters and their impact on the quantum signal. The quantum error corrector is responsible for applying QEC techniques to the quantum signal. The adaptive optics system is responsible for adjusting the quantum signal’s path in real time. The signal enhancer and recoverer are responsible for processing the quantum signal to improve the SNR and recover the original quantum state.
@@ IV-B Predicted Results and Analysis
Figure 3: Graph showcasing the predicted secure information transmission rates of the proposed transmission scheme, the secure coding based on JEEC Coding [15], the secure coding based on LPS codes, and Cs for a practical QSDC system, without the consideration of the loss caused by the delayed fiber [26]
Our estimations for our proposed system’s performance are substantiated by empirical evidence and mathematical proofs from recent research. Studies by Roffe [33] on quantum LDPC codes and by Ghilea [34] on quasi-cyclic multi-edge LDPC codes demonstrate significant advancements in error correction, decoder efficiency, and communication distance for quantum key distribution systems. Furthermore, Mele, Lami, and Giovannetti [35] Investigated how quantum communication technologies withstand noise and attenuation, laying a robust foundation for our forecasts. The results collectively highlight the capability of our suggested system to improve upcoming QSDC systems, providing robust empirical and mathematical backing for our calculations.
@@ V-A Technical Implementation of LF QSDC
For extensive free-space communication, the one-way transmission methods (RECON protocol and QKPC protocol) might offer more benefits than the two-way protocols, given the various elements in the free-space channel influencing communication [36]. The conveyance of data across extensive distances via a free-space channel is fraught with multiple difficulties. Factors such as beam-spreading, atmospheric turbulence, absorption and scattering, background light (sunlight), geometrical loss, and weather conditions could all affect communication [37, 38, 39]. Figure 4 summarizes the main characteristics of free-space channel and their impacts on optical signal transmission.
Figure 4: Summarizes the main characteristics of free-space channel and their impacts on optical signal transmission[36, 37]
@@ V-A2 Experimental Implementation
Attaining a low QBER in experiments is a key mea- sure of QSDC’s resilience and dependability in free- space channels [44, 45]. A minimal QBER indicates the ability to convey quantum data with great accuracy, crucial for ensuring secure communication [46, 47]. Attaining a data transfer speed of 500 bits per second across a 10-meter free-space channel showcases the pro- tocol’s capability for effective data transfer [43].
@@
Included in the experimental arrangement were sys- tems for identifying eavesdropping, an essential element of secure quantum communication. The protocol’s ca- pacity to adjust pulses for security verifications and se- cure information encoding showcases its effectiveness in maintaining communication confidentiality and integrity, even amidst possible risks [43].
@@ V-B1 Feasibility of Satellite Communication with LF QSDC
Earlier studies on the satellite-to-ground transfer of quantum states have yielded encouraging outcomes, with QBERs remaining within the tolerable limits for safe quantum communication[48, 49, 50]. The ability to transmit effectively across distances from 2,200 km to more than 36,000 km, maintaining link losses in the range of 100 to 110 dB, demonstrates that basic physics enables quantum communication across the extensive distances needed for satellite communication[50, 51].
The DL04 protocol, devoid of memory, plays a crucial role in LF QSDC by simplifying the processes of quantum state storage and management [19]. The DL04 protocol’s LF flexibility in adapting to free-space environments renders it an appealing choice for satellite communication, especially in addressing the difficulties of transmitting quantum signals across Earth’s atmosphere. The incorporation of quantum-aware LDPC coding into LF QSDC markedly enhances the system’s error rectification abilities, tackling the problems caused by extensive transmission and atmospheric influences [52, 53]. Furthermore, the use of PAT technologies helps sustain a consistent and precise communication connection between satellites and terrestrial stations [54].
@@ V-B2 Expanding Feasibility with LF QSDC
Enhancing the feasibility of LF QSDC involves comprehensive advancements across various technical and regulatory dimensions. Key initiatives include refining beam propagation and encoding methods to counter atmospheric disturbances and environmental disruptions as shown in Figure 5, ensuring high communication accuracy and low Quantum Bit Error Rate (QBER) across vast distances [55, 56]. Additionally, integrating quantum repeaters and relay satellites expands the scope of QSDC, enabling a global quantum communication network that overcomes spatial limitations inherent in direct quantum communication [57]. Improvements in daylight operation capabilities through advanced detectors and narrowband filters are crucial for the system’s effectiveness in satellite applications, addressing challenges posed by solar radiation [58].
@@
Additionally, LF QSDC contributes to Web 3.0’s goals of user privacy and anonymity by establishing secure, direct communication channels that eliminate the need for intermediaries in key exchanges or transaction verifications. This approach not only strengthens confidentiality within the network but also introduces an additional layer of transaction security based on quantum mechanics, offering a robust defense against potential quantum cryptographic attacks [21, 68].
@@ VI-A2 Distance
Yin explored free-space QSDC’s capability for long-distance communication through free-space channels and successfully demonstrated quantum teleportation over a 97-km channel and entanglement distribution over a 101.8-km two-link channel, achieving an average fidelity of 80.4% for six initial states despite significant channel loss (35-53 dB for teleportation and 66-85 dB for entanglement distribution)[74]. This showcases LF QSDC’s potential for satellite-based quantum communication, significantly extending the feasible communication distance beyond traditional QSDC methods. Achieving free-space QSDC over distances of 100 kilometers has been validated through other successful experiments in quantum teleportation and the distribution of entanglement across distances surpassing 100 km[75].

### v5

@@ I Introduction
In the evolution of satellite-based QKD, significant strides have been made since the launch of the Micius satellite in 2017, demonstrating secure quantum key transmission over intercontinental distances [3, 5, 6]. These advancements include ground-to-satellite quantum teleportation [3], air-to-ground quantum communication [4], and satellite-based entanglement distribution [5]. However, challenges such as substantial signal loss, atmospheric interference, and the need for precise alignment and tracking remain [4, 5, 6].
Zhou et al. addressed some of these issues with device-independent QSDC protocols [7]. Zhang et al. demonstrated QSDC over 100 km, highlighting its scalability [8]. Liu et al. discussed the infrastructure needed for global QSDC networks, including numerous satellites and ground stations [9]. Zhou et al. emphasized the importance of integrating quantum repeaters to extend QSDC’s reach [10]. And more works by many scientists continue to advance the field of QSDC[11, 12, 13]
Despite many advancements, practical deployment of satellite-based QSDC still faces significant hurdles, requiring continued research and development to achieve scalable, secure quantum communication networks[14].
Challenges with QKD for Global Scale Communication: QKD systems experience significant signal loss and attenuation over long distances, limiting practical distances without quantum repeaters to a few hundred kilometers. Atmospheric disturbances further exacerbate signal degradation in free-space QKD, making it highly susceptible to environmental factors. For instance, substantial signal loss and atmospheric interference hinder the efficiency of QKD in long-distance communications
Extending QKD reach beyond terrestrial limits with satellites introduces additional challenges, including precise satellite alignment and significant signal loss during transmission. Implementing a global QKD network requires substantial infrastructure, including numerous satellites and ground stations, and the integration of quantum repeaters. The high cost and complexity of this infrastructure pose barriers to widespread adoption
Advantages of QSDC for Global Scale Communication In contrast, QSDC protocols, particularly our solution LF QSDC, offer more practical solutions for global communication. QSDC allows direct transmission of secure messages without prior key exchange, reducing communication complexity and eliminating intermediate key management. Additionally, QSDC protocols incorporate eavesdropping detection through quantum state disturbance, ensuring real-time detection of interference. Technologies like Quantum-Aware Low-Density Parity-Check (LDPC) coding, Pointing, Acquisition, and Tracking (PAT) systems, and Atmospheric Quantum Correction Algorithms (AQCA) improve LF QSDC reliability over long distances. These technologies address key challenges in satellite-based free-space communication, ensuring precise alignment and minimizing signal loss due to atmospheric disturbances. LF QSDC’s reliance on advanced error correction and adaptive technologies reduces the need for extensive infrastructure compared to QKD, making it more scalable and cost-effective. By leveraging existing satellite and communication infrastructure with minimal modifications, LF QSDC can be integrated seamlessly into current systems.
@@
Recent advancements in QSDC and supporting fields have shown significant theoretical and experimental progress, enhancing the feasibility of direct quantum communications[16, 17, 18, 19, 20, 21]. Noteworthy developments include the realization of QSDC over 100 km using time-bin and phase quantum states underlines the scalability of these technologies[10].
@@
The memory-free DL04 protocol that FL QSDC employs reduces the complexity of quantum state storage and management, making the system more practical and robust for real-world applications [15] as shown in Table I. Our novel 1uantum-aware LDPC coding scheme, PAT technologies, and AQCA mitigate issues related to atmospheric turbulence and alignment errors, which are prominent in satellite-based free-space quantum communication. Our innovative features are based on the developments of these works [22, 23, 24, 25, 26, 27].
@@ II-B Measurement Protocols and Two-way Communication
After measurement, Bob announces his choice of $\theta$ and the results of his measurements. Alice and Bob then jointly analyze the outcomes to check if $\theta$ aligns with $a_{2}$ ($a_{2}\oplus\theta=0$ or $1$), ensuring the correct basis was used and the transmission was secure. This process embodies the two-way communication required by the DL04 protocol, where feedback from Bob influences potential retransmissions by Alice or adjustments in the encoding strategy to correct any discrepancies or enhance security.
@@ III-A3 Security Enhancement Techniques
  • The estimation can be formulated as solving a set of linear inequalities derived from decoy state intensities and detection rates.
@@ III-C1 Design
The AQCA consists of four main modules: a disturbance modeler, a quantum error corrector, an adaptive optics system, and a signal enhancer and recoverer. The disturbance modeler is responsible for estimating the atmospheric parameters and their impact on the quantum signal. The quantum error corrector is responsible for applying QEC techniques to the quantum signal. The adaptive optics system is responsible for adjusting the quantum signal’s path in real-time. The signal enhancer and recoverer are responsible for processing the quantum signal to improve the SNR and recover the original quantum state.
@@ IV-B Predicted Results and Analysis
Figure 3: Graph showcasing the predicted secure information transmission rates of the proposed transmission scheme, the secure coding based on JEEC Coding, the secure coding based on LPS codes, and Cs for a practical QSDC system, without the consideration of the loss caused by the delayed fiber
Our estimations for our proposed system’s performance are substantiated by empirical evidence and mathematical proofs from recent research such as works by Hu et al. and Babar et al. and[30, 31]. Specifically, studies by Roffe [33] on quantum LDPC codes and by Ghilea [34] on quasi-cyclic multi-edge LDPC codes demonstrate significant advancements in error correction, decoder efficiency, and communication distance for quantum key distribution systems. Furthermore, Mele, Lami, and Giovannetti [35] Investigated how quantum communication technologies withstand noise and attenuation, laying a robust foundation for our forecasts. The results collectively highlight the capability of our suggested system to improve upcoming QSDC systems, providing robust empirical and mathematical backing for our calculations.
@@ V-A Technical Implementation of LF QSDC
For extensive free-space communication, the one-way transmission methods (RECON protocol and QKPC protocol) might offer more benefits than the two-way protocols, given the various elements in the free-space channel influencing communication [36]. The conveyance of data across extensive distances via a free-space channel is fraught with multiple difficulties. Factors such as beam-spreading, atmospheric turbulence, absorption and scattering, background light (sunlight), geometrical loss, and weather conditions could all affect communication [37, 38, 39]. Figure 4 summarizes the main characteristics of the free-space channel and their impacts on optical signal transmission.
Figure 4: Summarizes the main characteristics of free-space channel and their impacts on optical signal transmission[62]
@@ V-A2 Experimental Implementation
Attaining a low QBER in experiments is a key measure of QSDC’s resilience and dependability in free-space channels [44]. A minimal QBER indicates the ability to convey quantum data with great accuracy, crucial for ensuring secure communication demonstrated by both works in QSDC and QKD [45, 46, 47]. Attaining a data transfer speed of 500 bits per second across a 10-meter free-space channel showcases the protocol’s capability for effective data transfer [62].
@@
Included in the experimental arrangement were also systems for identifying eavesdropping, an essential element of secure quantum communication. A protocol’s capacity to adjust pulses for security verifications and secure information encoding showcases its effectiveness in maintaining communication confidentiality and integrity is crucial and achieveavle, even amidst possible risks as demonstrated by this experiment and Pan et al.[43].
@@ V-B1 Feasibility of Satellite Communication with LF QSDC
Earlier studies on the satellite-to-ground transfer of quantum states have yielded encouraging outcomes, with QBERs remaining within the tolerable limits for safe quantum communication[48, 49, 50]. The ability to transmit effectively across distances from 2,200 km to more than 36,000 km, maintaining link losses in the range of 100 to 110 dB, demonstrates that basic physics enables quantum communication across the extensive distances needed for satellite communication[50].
The DL04 protocol, devoid of memory, plays a crucial role in LF QSDC by simplifying the processes of quantum state storage and management [19]. The DL04 protocol’s LF flexibility in adapting to free-space environments renders it an appealing choice for satellite communication, especially in addressing the difficulties of transmitting quantum signals across Earth’s atmosphere. The incorporation of quantum-aware LDPC coding into LF QSDC markedly enhances the system’s error rectification abilities, tackling the problems caused by extensive transmission and atmospheric influences [51, 52, 53]. Furthermore, the use of PAT technologies helps sustain a consistent and precise communication connection between satellites and terrestrial stations [54].
@@ V-B2 Expanding Feasibility with LF QSDC
Enhancing the feasibility of LF QSDC involves comprehensive advancements across various technical and regulatory dimensions. Key initiatives include refining beam propagation and encoding methods to counter atmospheric disturbances and environmental disruptions as shown in Figure 5, ensuring high communication accuracy and low Quantum Bit Error Rate (QBER) across vast distances [55, 56]. Additionally, integrating quantum repeaters and relay satellites expands the scope of QSDC, enabling a global quantum communication network that overcomes spatial limitations inherent in direct quantum communication [57]. Improvements in daylight operation capabilities through advanced detectors and narrowband filters are crucial for QSDC’s effectiveness in satellite applications, addressing challenges posed by solar radiation [58].
@@
Additionally, LF QSDC contributes to Web 3.0’s goals of user privacy and anonymity by establishing secure, direct communication channels that eliminate the need for intermediaries in key exchanges or transaction verifications. This approach not only strengthens confidentiality within the network but also introduces an additional layer of transaction security based on quantum mechanics, offering a robust defense against potential quantum cryptographic attacks [68].
@@ VI-A2 Distance
Yin explored free-space QSDC’s capability for long-distance communication through free-space channels and successfully demonstrated quantum teleportation over a 97-km channel and entanglement distribution over a 101.8-km two-link channel, achieving an average fidelity of 80.4% for six initial states despite significant channel loss (35-53 dB for teleportation and 66-85 dB for entanglement distribution)[74]. This showcases LF QSDC’s potential for satellite-based quantum communication, significantly extending the feasible communication distance beyond traditional QSDC methods. Achieving free-space QSDC over distances of 100 kilometers has been validated through other successful experiments in quantum teleportation and the distribution of entanglement across distances surpassing 100 km[8, 75].